Given f(x) = x2 + x – 9, find f(0).
Answer: f(0) =

Answer by rapaljer(4671)

Given f(x) = x^2 + x – 9, find f(0).

f(x) = x^2 + x – 9
f(0)= o^2 + 0 - 9
f(0) = -9

R^2 at SCC